# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v2

@@ 2.2 Autonomous flying inside forests
Even though only a few studies have yet utilized robotic under-canopy drones in DBH estimation, drone autonomy itself has been an actively researched topic in the robotics community during the past few years. Recent advances in mobile robotics and the enhanced computational capabilities of small embedded computers have made it realistic to build drones capable of real-time trajectory planning in GNSS-denied cluttered environments. Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments. Most importantly, the authors of these methods have also shared their source code to be openly available for the research community to encourage more development in the field. While the methods utilize various sensor setups, they can be divided into LiDAR-based methods (Liu et al. 2022; Prabhu et al. 2024; Ren et al. 2022; Liu, Ren, and Zhang 2024; Ahmad et al. 2022) and camera-based methods (Campos-Macías et al. 2021; Zhou et al. 2022; Loquercio et al. 2021; Nguyen et al. 2023) based on the sensor utilized for the sensing of the obstacles in the flight path. Since GNSS is not a suitable method for localization under the forest canopy, these methods rely on various Simultaneous localization and mapping (SLAM) or odometry algorithms and onboard sensor data to obtain the pose estimate in a local coordinate system. Most of these methods tackle the problem of collision-free trajectory planning using 3D grid maps, but the solutions by Loquercio et al. (2021) and Nguyen et al. (2023) replaced individual mapping and trajectory planning subtasks by deep neural networks mapping the depth images directly to collision-free trajectories. In addition, the methods proposed by Zhou et al. (2022) and Ahmad et al. (2022) concentrated a swarm of drones instead of autonomous flying of an individual drone.
@@
However, most of these methods are typically validated by a single successful autonomous test flight in a specific forest environment. Only four of the articles reported multiple test flights inside real forests (Loquercio et al. 2021; Ren et al. 2022; Liu, Ren, and Zhang 2024; Prabhu et al. 2024). From these, (Loquercio et al. 2021) and (Prabhu et al. 2024) were the only methods tested at least in two different forest types. Moreover, since methods by (Loquercio et al. 2021), (Ren et al. 2022), and (Liu, Ren, and Zhang 2024) were all targeting high-speed flight, the test forests reported in those studies were relatively sparse, with a low amount of low-hanging branches and understory vegetation. The autonomous trajectory planning method by Prabhu et al. (2024) was also tested only in environments where the space between the trees was larger than in typical forests, and the data collecting flights in denser forests were manually piloted.
@@
Therefore, open questions remain regarding the applicability and reliability of these algorithms in diverse forest environments, especially in challenging boreal forests.
@@ 3.1 Autonomous navigation algorithms
In this study, the objective was to develop a prototype lightweight camera-based drone capable of autonomously avoiding obstacles under the forest canopy and operating without GNSS. The drone was required to fly autonomously from the takeoff location to a specified goal point while recording stereo-camera data during the flight. This prototype will later serve as a development platform for more extended and sophisticated data collection tasks.
Among the open-source solutions presented in Section 2.2, the solution by Zhou et al. (2022) was chosen as the base of the prototype. The solution was the only open-sourced camera-based solution proven to work with only one stereo camera making it possible to build a smaller and lighter platform. Additionally, the test environment reported in the original article (dense bamboo forest) was the most difficult among the proposed solutions, even though it had a highly different structure than the typical boreal forests.
@@ 3.1.1 EGO-Planner-v2
The source code of the planner algorithm by Zhou et al. (2022) was named in Github as EGO-Planner-v2 (ZJU FAST Lab Team 2022a). Although EGO-Planner-v2 is designed for swarm flying, in this study, the algorithm was modified for a single drone system to focus on navigation performance in challenging forest environments. The planner runs on top of the Robot Operating System (ROS) (Quigley et al. 2009). The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (Moravec and Elfes 1985). The mapping module uses depth images as an input, and the system adopts the fixed-sized circular buffers proposed by Usenko et al. (2017) for maintaining the local map. The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map, respectively. The virtual floor and the virtual ceiling restrict the altitude at which the algorithm can plan the paths.
In the trajectory planning, EGO-Planner-v2 uses a MINCO (minimum control) (Wang et al. 2022) trajectory representation designed for differentially flat systems, such as quadrotors (Mellinger and Kumar 2011). Planned trajectories are represented as piece-wise polynomial splines with decoupled temporal and spatial parameters (Zhou et al. 2022). The trajectories are optimized iteratively by minimizing the weighted sum of metrics defined for trajectory smoothness and flying time. The feasibility of the planned trajectory is forced by deforming the shape to avoid obstacles and by restricting the magnitudes of trajectory velocity, acceleration, and jerk to fulfill the dynamical constraints of the drone. For computational efficiency, the trajectory constraints are enforced via integrals of penalty functions with large penalty weights, and the integrals are evaluated by a finite sum of equally spaced samples along the timeline. In a simplified form, the optimization of time-dependent objectives $J$ in one piece of the trajectory can be represented by
@@ 3.1.2 VINS-Fusion
VINS-Fusion supports loop detection-based relocalization and global pose graph optimization to reduce the accumulated drift in the pose estimate. The loops are detected by storing additional corner features as BRIEF descriptors (Calonder et al. 2010) to DBoW2 (Galvez-López and Tardos 2012) bag-of-word database. After a loop detection, the relocalization process aligns the latest window to past poses and optimizes it utilizing also the feature correspondences with past poses. In addition to sliding window relocalization, the global pose graph is also optimized after loop detections to enforce the global consistency of the poses. Since the roll and pitch angle estimates are drift accumulation-free due to full observability from the gravity measurement of the IMU, the pose graph optimization is performed in 4 degrees of freedom.
@@
The parameters used in the configuration of VINS-Fusion were chosen so that real-time performance was guaranteed. In the initial tests, the IMU noise values used during the calibration were over-optimistic for VINS-Fusion causing instability to the estimates, so the IMU noise values were increased more. All VINS-Fusion configuration parameters are listed in Table 2.
@@ 3.1.3 Tracking controller
The original EGO-Planner-v2 article (Zhou et al. 2022) did not specify the utilized method for trajectory tracking. In this study, an open-source tracking controller (ZJU FAST Lab Team 2022b) by Zhejiang University was used. The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner and converts those to attitude and thrust setpoints. The setpoints are forwarded to the PX4 flight controller software (PX4 Community 2023) responsible for the low-level attitude and angular rate controls.
@@ 3.2 Drone hardware
RealSense D435 is based on active stereoscopy, and the camera has an infrared laser projector emitting dots that can be used for triangulation by the stereo camera. This can improve the triangulation in low-light conditions and to homogeneous surfaces that contain no features. However, the dots of the laser emitter cause problems for VIO since they are moving with the camera. In the previous set-up described by (Karjalainen et al. 2023), a modified version of the ROS-RealSense driver by Zhou et al. (2021b) was used. The modified driver turns the laser emitter of the camera constantly on and off, making it possible to produce high-quality depth images in every other frame and stereo images without the laser dots in every other frame. However, this solution comes with the drawback of halving the output frame rate of the camera.
In the first test flights described by (Karjalainen et al. 2023), difficulties were encountered in detecting small branches with a depth camera resolution of $640\times 480$, leading to the decision to increase the depth camera resolution to a maximum value of $1280\times 720$. The stereo images for VIO were downsampled at a resolution of $640\times 360$ to ensure real-time performance. As the maximum frame rate of D435 at a resolution of $1280\times 720$ is 30 Hz, the modified driver version would have resulted in a low frame rate of 15 Hz. To ensure the VIO performance in quick turns, the frame rate of 30 Hz was preferred, and hence, the IR emitter was disabled, making the camera function as a passive IR stereo camera. The choice can be further justified with the fact that the usefulness of the IR emitter in forest conditions is questionable since sunlight contains infrared light that can interfere with the emitted artificial features(Kuan, Ee, and Wei 2019), and the forests typically do not contain large homogeneous surfaces. The Realsense D435 onboard camera was configured to use the built-in automatic exposure control.
The camera-IMU system was calibrated with an open-source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012; Furgale, Rehder, and Siegwart 2013; Maye, Furgale, and Siegwart 2013). In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were calculated from recorded data where Aprilgrid (Olson 2011) was used as a calibration target. The IMU noise parameters needed for the calibration were first obtained with the open source tool Allan Variance ROS (Oxford Dynamic Robot Systems Group 2022), and the values obtained were multiplied by 10 before the calibration as suggested in the wiki of the kalibr project (ETH Zürich Autonomous Systems Lab 2023).
@@ 3.3 Test sites
For reference purposes, circular GCPs provided by Agisoft Metashape (Agisoft LLC, St. Petersburg, Russia) were placed within the forest plots used in the flight tests. The targets had a diameter of 24.7 cm and the middle dots had a size of 7 cm. Relative distances between the GCPs were measured with Leica Nova TS60 total station with a precision of approximately 2 mm (Leica Geosystems 2020). Figure 5 shows GCPs in ’Evo-difficult’ during the test flights.
@@
GCPs were not used in the third test forest, ’Paloheinä’. The experiment there aimed to test the loop detection performance of VINS-Fusion in a homogeneous forest environment without any artificial targets. The data were collected by walking with the drone. The estimated average tree density in the area was approximately 2000 trees/ha.
@@ 3.4 Flight tests
Test flights in Evo were conducted on 18th October 2023. Multiple test flights were performed to evaluate the forest-type effects on the performance of autonomous navigation. In total, seven test flights were conducted in the ’Evo-medium’ test area and nine test flights in the ’Evo-difficult’ test area. In each test environment, either the takeoff location or the takeoff heading was altered during the test flight campaign to force the drone to follow different paths. In the ’Evo-medium’ test area, the takeoff location was changed after five test flights. In the ’Evo-difficult’ test area, the takeoff heading was changed after six test flights. The target flight velocity was 1 m/s in all flight tests. The goal was set 36 m forward in the first takeoff location in the ’Evo-medium’ test area, 34 m forward in the second takeoff location in the ’Evo-medium’ test area, and 42 m forward in the takeoff location in the ’Evo-difficult’ test area. The maximum flight height was set to 2.25 m relative to the takeoff location in the ’Evo-medium’ test area. For the first test flight in the ’Evo-difficult’ test area, the maximum flight height was set to 2.5 m with respect to the takeoff location and was increased to 2.75 m for the rest of the flights.
The campaign to evaluate the loop detection capability took place on 26th January 2024. Stereo images and IMU data were collected during two walks inside the forest while carrying the drone by hand. The paths of the walks were approximately 340 m and 240 m long, each containing multiple loops where the same scene was observed multiple times. The data collection took place during winter, with the forest covered in snow.
@@ 3.5 Structure-from-Motion Multi-View Stereo based reconstruction
Stereo image processing was carried out using Agisoft Metashape software (Agisoft LLC, St. Petersburg, Russia) (Metashape Development Team 2021) following similar processing steps as introduced by Viljanen et al. (2018). Every sixth stereo pair was used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m. In the image orientation processing, the quality setting was set to high meaning that two times downsampled images were used. The settings for the number of key points and tie points per image were 40,000 and 4,000, respectively. To enhance the quality of the post-processed reference trajectories, GCPs were used in the processing with an accuracy setting of ±$\pm 0.005$ m. Automatic outlier removal was performed using gradual selection tools of the software based on re-projection error, reconstruction uncertainty, and projection accuracy. Subsequently, a bundle adjustment was carried out, involving optimization of the sparse point cloud along with the interior (IOP) and the exterior orientation parameters (EOP). Following that, a dense point cloud was calculated using the highest quality setting and mild depth filtering. Finally, the EOPs and dense point clouds were exported.
@@ 3.6 Tree parameter estimation
The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments, and 5-s-long temporal segments. The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus, the temporal segmentation was not performed for the TLS data. The points potentially corresponding to the tree stems were detected by clustering the points with density-based clustering (DBSCAN, (Ester et al. 1996)) and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle. The accepted clusters were further grouped into tree stems by applying DBSCAN for the center points of the clusters, and the growth direction of each tree was estimated with principal component analysis (PCA). The points within accepted clusters were projected into the plane perpendicular to the growth direction for the final diameter estimation. The diameters were processed into a stem curve by automatically removing clear outliers, followed by fitting a smoothing spline to the remaining diameter estimates as a function of height. Finally, the DBH estimates were calculated by interpolating the stem curve at the height of 1.3 m above the ground. For a detailed description of the method, the reader is referred to Hyyppä et al. (2020b, a).

### v3

@@ 2.2 Autonomous flying inside forests
Even though only a few studies have yet utilized robotic under-canopy drones in DBH estimation, drone autonomy itself has been an actively researched topic in the robotics community during the past few years. Recent advances in mobile robotics and the improved computational capabilities of small embedded computers have made it realistic to build drones capable of real-time trajectory planning in cluttered GNSS-denied environments. Although technology is still in the early stages of development, some recent studies have proposed various methods that have been successfully tested in forest environments. Most importantly, the authors of these methods have also shared their source codes to make them openly available to the research community to encourage more development in the field. Although the methods utilize various sensor setups, they can be divided into LiDAR-based methods (Liu et al. 2022; Prabhu et al. 2024; Ren et al. 2022; Liu, Ren, and Zhang 2024; Ahmad et al. 2022) and camera-based methods (Campos-Macías et al. 2021; Zhou et al. 2022; Loquercio et al. 2021; Nguyen et al. 2023) based on the sensor utilized for the detection of obstacles in the flight path. Since GNSS is not a suitable method for localization under the forest canopy, these methods rely on various Simultaneous localization and mapping (SLAM) or odometry algorithms and onboard sensor data to obtain the pose estimate in a local coordinate system. Most of these methods tackle the problem of collision-free trajectory planning using 3D grid maps, but the solutions by Loquercio et al. (2021) and Nguyen et al. (2023) replaced individual mapping and trajectory planning subtasks by deep neural networks mapping the depth images directly to collision-free trajectories. In addition, the methods proposed by Zhou et al. (2022) and Ahmad et al. (2022) concentrated a swarm of drones instead of the autonomous flying of an individual drone.
@@
However, most of these methods were validated by a single successful autonomous test flight in a specific forest environment. Only four of the articles reported multiple test flights inside real forests (Loquercio et al. 2021; Ren et al. 2022; Liu, Ren, and Zhang 2024; Prabhu et al. 2024). From these, (Loquercio et al. 2021) and (Prabhu et al. 2024) were the only methods tested at least in two different forest types. Furthermore, since the methods by (Loquercio et al. 2021), (Ren et al. 2022), and (Liu, Ren, and Zhang 2024) were all targeting high-speed flight, the test forests reported in these studies were relatively sparse, with a low amount of low-hanging branches and understory vegetation. The autonomous trajectory planning method by Prabhu et al. (2024) was also tested only in environments where the space between the trees was larger than in typical forests, and the data-collecting flights in denser forests were manually piloted.
@@
Therefore, open questions remain regarding the applicability and reliability of these algorithms in diverse forest environments, especially challenging boreal forests.
@@ 3.1 Autonomous navigation algorithms
In this study, the objective was to develop a prototype lightweight camera-based drone capable of avoiding obstacles autonomously under the forest canopy and operating without GNSS. The drone was required to fly autonomously from the takeoff location to a specified goal point while recording stereo camera data during the flight. This prototype will later serve as a development platform for more extended and sophisticated data collection tasks.
Among the open source solutions presented in Section 2.2, the solution by Zhou et al. (2022) was chosen as the basis of the prototype. The solution was the only open source camera-based solution that was proven to work with only one stereo camera, making it possible to build a smaller and lighter platform. Furthermore, the test environment reported in the original article (dense bamboo forest) was the most difficult among the proposed solutions, although it had a highly different structure from typical boreal forests.
@@ 3.1.1 EGO-Planner-v2
The source code of the planner algorithm by Zhou et al. (2022) was named on GitHub as EGO-Planner-v2 (ZJU FAST Lab Team 2022a). Although EGO-Planner-v2 is designed for swarm flying, in this study, the algorithm was modified for a single drone system to focus on navigation performance in challenging forest environments. The planner runs on top of the Robot Operating System (ROS) (Quigley et al. 2009). The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (Moravec and Elfes 1985). The mapping module uses depth images as input and the system adopts the fixed-size circular buffers proposed by Usenko et al. (2017) to maintain the local map. The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map, respectively. The virtual floor and the virtual ceiling restrict the altitude at which the algorithm can plan the paths.
In trajectory planning, EGO-Planner-v2 uses a MINCO (minimum control) (Wang et al. 2022) trajectory representation designed for differentially flat systems, such as quadrotors (Mellinger and Kumar 2011). The planned trajectories are represented as piece-wise polynomial splines with decoupled temporal and spatial parameters (Zhou et al. 2022). The trajectories are optimized iteratively by minimizing the weighted sum of metrics defined for trajectory smoothness and flying time. The feasibility of the planned trajectory is forced by deforming the shape to avoid obstacles and by restricting the magnitudes of the trajectory velocity, acceleration, and jerk to meet the dynamical constraints of the drone. For computational efficiency, trajectory constraints are enforced through integrals of penalty functions with large penalty weights, and the integrals are evaluated by a finite sum of equally spaced samples along the timeline. In a simplified form, the optimization of time-dependent objectives $J$ in one piece of the trajectory can be represented by
@@ 3.1.2 VINS-Fusion
VINS-Fusion supports loop detection-based relocalization and global pose graph optimization to reduce the accumulated drift in the pose estimate. The loops are detected by storing additional corner features as BRIEF descriptors (Calonder et al. 2010) to a DBoW2 (Galvez-López and Tardos 2012) bag-of-word database. After a loop detection, the relocalization process aligns the latest window to past poses and optimizes it utilizing also the feature correspondences with past poses. In addition to sliding-window relocalization, the global pose graph is also optimized after loop detections to enforce the global consistency of the poses. Since the roll and pitch angle estimates are drift accumulation-free due to full observability from the gravity measurement of the IMU, the pose graph optimization is performed in 4 degrees of freedom.
@@
The parameters used in the configuration of VINS-Fusion were chosen so that real-time performance was guaranteed. In the initial tests, the IMU noise values used during the calibration were over-optimistic for VINS-Fusion, causing instability to the estimates, so the IMU noise values were increased further. All VINS-Fusion configuration parameters are listed in Table 2.
@@ 3.1.3 Tracking controller
The original EGO-Planner-v2 article (Zhou et al. 2022) did not specify the method for trajectory tracking. In this study, an open source tracking controller (ZJU FAST Lab Team 2022b) by Zhejiang University was used. The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner and converts those to attitude and thrust setpoints. The setpoints are forwarded to the PX4 flight controller software (PX4 Community 2023) responsible for the low-level attitude and angular rate controls.
@@ 3.2 Drone hardware
The RealSense D435 onboard camera was configured to the maximum image resolution ($1280\times 720$) and the update frequency (30 Hz). The exposure value was set to use the built-in automatic exposure control. The stereo images for VIO were downsampled at a resolution of $640\times 360$ to ensure real-time performance.
Stereoscopic RealSense D435 has an infrared laser projector emitting dots, improving the triangulation in low light conditions and for homogeneous surfaces. However, the dots of the laser emitter cause problems for VIO since they are moving with the camera. Therefore, in this study, the infrared emitter was disabled. Since sunlight contains infrared light, which interferes with the artificial features emitted (Kuan, Ee, and Wei 2019), and forests typically do not contain homogeneous surfaces, disabling the emitter should not affect the camera performance in outdoor conditions.
The camera-IMU system was calibrated with an open source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012; Furgale, Rehder, and Siegwart 2013; Maye, Furgale, and Siegwart 2013). In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were estimated from a recorded data sequence, where Aprilgrid (Olson 2011) was used as a calibration target. The IMU noise parameters needed for the calibration were first obtained with the open source tool Allan Variance ROS (Oxford Dynamic Robot Systems Group 2022), and the values obtained were multiplied by 10 before the calibration as suggested in the wiki of the kalibr project (ETH Zürich Autonomous Systems Lab 2023).
@@ 3.3 Test sites
For reference purposes, circular GCPs provided by Agisoft Metashape (Agisoft LLC, St. Petersburg, Russia) were placed within the forest plots used in the flight tests. The targets had a diameter of 24.7 cm and the middle dots had a diameter of 7 cm. Relative distances between the GCPs were measured with a Leica Nova TS60 total station with a precision of approximately 2 mm (Leica Geosystems 2020). Figure 5 shows the GCPs in ’Evo-difficult’ during the test flights.
@@
GCPs were not used in the third test forest, ’Paloheinä’. The experiment there aimed to test the loop detection performance of VINS-Fusion in a homogeneous forest environment without artificial targets. The data sequences were collected by walking with the drone. The estimated average tree density in the area was approximately 2000 trees/ha.
@@ 3.4 Flight tests
Test flights in Evo were conducted on 18th October 2023. Multiple test flights were conducted to evaluate forest-type effects on autonomous navigation performance. In total, seven test flights were conducted in the ’Evo-medium’ test area and nine test flights in the ’Evo-difficult’ test area. In each test environment, either the takeoff location or the takeoff heading was altered during the test flight campaign to force the drone to follow different paths. In the ’Evo-medium’ test area, the takeoff location was changed after five test flights. In the ’Evo-difficult’ test area, the takeoff heading was changed after six test flights. The target flight velocity was 1 m/s in all flight tests. The goal was set 36 m forward in the first takeoff location in the ’Evo-medium’ test area, 34 m forward in the second takeoff location in the ’Evo-medium’ test area, and 42 m forward in the takeoff location in the ’Evo-difficult’ test area. The maximum flight height was set to 2.25 m relative to the takeoff location in the ’Evo-medium’ test area. For the first test flight in the ’Evo-difficult’ test area, the maximum flight height was set to 2.5 m with respect to the takeoff location and was increased to 2.75 m for the rest of the flights.
The campaign to evaluate the loop detection capability took place on 26th January 2024. Stereo images and IMU data were collected during two walks inside the forest while carrying the drone by hand. The paths of the walks were approximately 340 m long and 240 m long, each containing multiple loops where the same scene was observed multiple times. Data collection was carried out during winter, with the forest covered in snow.
@@ 3.5 Structure-from-Motion Multi-View Stereo based reconstruction
Stereo image processing was carried out using Agisoft Metashape software (Agisoft LLC, St. Petersburg, Russia) (Metashape Development Team 2021) following similar processing steps as introduced by Viljanen et al. (2018). Every sixth stereo pair was used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m. In the image orientation processing, the quality setting was set to high, meaning that two times downsampled images were used. The settings for the number of key points and tie points per image were 40,000 and 4,000, respectively. To enhance the quality of the post-processed reference trajectories, GCPs were used in the processing with an accuracy setting of $\pm 0.005$ m. Automatic outlier removal was performed using the gradual selection tools of the software based on re-projection error, reconstruction uncertainty, and projection accuracy. Subsequently, a bundle adjustment was carried out, involving optimization of the sparse point cloud along with the interior (IOP) and the exterior orientation parameters (EOP). Following that, a dense point cloud was calculated using the highest quality setting and mild depth filtering. Finally, the EOPs and dense point clouds were exported.
@@ 3.6 Tree parameter estimation
The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments and 5-s-long temporal segments. The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus, the temporal segmentation was not performed for the TLS data. The points potentially corresponding to the tree stems were detected by clustering the points with density-based clustering (DBSCAN, (Ester et al. 1996)) and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle. The accepted clusters were further grouped into tree stems by applying DBSCAN for the center points of the clusters, and the growth direction of each tree was estimated with principal component analysis (PCA). The points within accepted clusters were projected into the plane perpendicular to the growth direction for the final diameter estimation. The diameters were processed into a stem curve by automatically removing clear outliers, followed by fitting a smoothing spline to the remaining diameter estimates as a function of height. Finally, the DBH estimates were calculated by interpolating the stem curve at a height of 1.3 m above the ground. For a detailed description of the method, the reader is referred to Hyyppä et al. (2020b, a).